If N drops of same size each having the same charge, coalesce to form a bigger drop. How will the following vary with respect to single small drop?              (3) 
 
(i) Total charge on bigger drop (ii) Potential on the bigger drop (iii) Capacitance 

Grade:12

2 Answers

Vikas TU
7 years ago
Volume would be conserved.
Hence,
V =nv
4/3*pi * R^3 = n*4/3*pi*r^3
R^3 = n*r^3...........(1)
And as charge is a scalar quantity,
Q = nq......(2)
From both eqns. we get,
Volt = KQ/R => Knq/(rn^(1/3))
Capicatance = Q/Volt => nq/[Knq/(rn^(1/3))] => r*n^(1/3)/K Farad. 
Tanush
7 years ago
 
Volume would be conserved.
Hence,
V =nv
4/3*pi * R^3 = n*4/3*pi*r^3
R^3 = n*r^3...........(1)
And as charge is a scalar quantity,
Q = nq......(2)
From both eqns. we get,
Volt = KQ/R => Knq/(rn^(1/3))
Capicatance = Q/Volt => nq/[Knq/(rn^(1/3))] => r*n^(1/3)/K Farad.